OpenStudy (anonymous):

Please help. urgent. Simplify the given expression: sqrt -64 over (7-6i)-(2-2i)

OpenStudy (anonymous):

\[i ^{2}=-1\] \[\sqrt{-64}=\sqrt{64 i ^{2}}\] Let A=7-6i and B=2-2i \[1=(A+B)\div(A+B)\] \[(8i \div(A-B))\times1=(8i \div(A-B))\times((A+B)\div(A+B))\] The resulted formula after simplification will be in the form of:\[(8i \times (A+B))/(A ^{2}-B ^{2})\] Just replace A and B with 7-6i and 2-2i, you'll get your answer. ;)